# Leveraging Text-to-Image Generation for Handling Spurious Correlation

### v1

@@ Abstract
Deep neural networks trained with Empirical Risk Minimization (ERM) perform well when both training and test data come from the same domain, but they often fail to generalize to out-of-distribution samples. In image classification, these models may rely on spurious correlations that often exist between labels and irrelevant features of images, making predictions unreliable when those features do not exist. We propose a technique to generate training samples with text-to-image (T2I) diffusion models for addressing the spurious correlation problem. First, we compute the best describing token for the visual features pertaining to the causal components of samples by a textual inversion mechanism. Then, leveraging a language segmentation method and a diffusion model, we generate new samples by combining the causal component with the elements from other classes. We also meticulously prune the generated samples based on the prediction probabilities and attribution scores of the ERM model to ensure their correct composition for our objective. Finally, we retrain the ERM model on our augmented dataset. This process reduces the model’s reliance on spurious correlations by learning from carefully crafted samples for in which this correlation does not exist. Our experiments show that across different benchmarks, our technique achieves better worst-group accuracy than the existing state-of-the-art methods.
@@ 1 Introduction
Inline with the compositional view of images [19], our method leverages textual inversion [6] and identifies the optimal tokens for the causal parts of each class. It then employs language-based segmentation [13] to identify the visual causal counterparts in the images. These parts are later employed by our technique for diffusion-based image generation that can balance the data. In Fig. 1, we show representative examples of the generated minority group samples computed by our technique by altering the causal parts of the majority class images from the training data. A key challenge in generating such samples is ensuring that they truly capture the desired properties for addressing the spurious correlation problem. Otherwise, the impact of balancing can be detrimental. Our processing pipeline leverages multiple components, and under-performance by any of those can potentially lead to a low quality image. To handle that, we introduce an effective pruning strategy based on the attribution score and the prediction probability distribution of the ERM model. Our approach evaluates how much a generated sample influences the model’s predictions, ensuring that only samples with a strong causal impact are retained. Moreover, by further leveraging attribution scores, our approach ensures higher-quality relevant images, leading to improved group balancing and better generalization.
@@ 3 Problem setup
Given a dataset $D_{tr}=\{(x_{i},y_{i})\}_{i=0}^{N}$, a typical classifier $f$ is aimed at estimating $P(y|x)$. We consider the compositional nature [30] of the samples $\{x_{i}\}_{i=0}^{N}$, where any sample consists of two types of features: core features ($c$) and spurious features ($s$). Core features (c) causally determine the correct label of the sample, while the spurious features (s) correlate with the label mainly due to irrelevant peculiarities of the dataset, e.g., class imbalances. These features are not inherently predictive. In other words, $P(y|s)\neq P(y)$.
@@
Based on different combinations of $(c,s)$, a dataset can be partitioned into distinct groups. The majority group comprises most of the samples in the dataset. Those samples also contain spurious feature ($s_{maj}$). Due to the correlation between $s_{maj}$ and the label, the classifier learns to rely on them for prediction. However, on encountering samples where this correlation does not hold, the classifier fails to generalize [29, 7]. We follow the common convention of referring to the data group on which $f$’s performance is the worst as the worst group.
@@ 4 Method
We introduce a three-stage approach designed to balance dataset groups effectively - see Fig. 2. Our first stage focuses on generating new data for minority groups using a diffusion-based generative model [22]. In the second stage, the generated data is pruned based on empirical risk minimization (ERM) loss and attribution scores, ensuring that only high-quality and relevant samples are retained. Finally, in the third stage, the ERM model is retrained on the newly balanced dataset, improving its robustness and mitigating spurious correlations. Below, we discuss these stages in detail.
Figure 2: An overview of our method - Dog/Cat classes used for illustration. (a) First, the token $C_{dog}$ is optimized by reconstructing samples from the Dog class using prompts that include the trainable $C_{dog}$. Then, for each batch in $\mathcal{D}_{cat}$, masks $M$ are generated using the LangSAM [13] with the prompt Class Name = ‘animal’. These masks, along with the input images and a textual prompt, are fed into the inpainting model to generate new images. The token $C_{dog}$ is incorporated to generate the causal features of the Dog class. (b) The generated samples are then passed through the ERM model and the integrated gradients (IG) module to compute relevant scores for filtering undesired generated images. Algorithm for this stage is detailed in Alg. 1. (c) The final set of generated samples is used to retrain the ERM model.
@@ 4.1 Generating data
Causal feature learning. Trained on enormous datasets, contemporary text-to-image (T2I) generative models can provide strong priors for natural image data augmentation [4]. We leverage this fact to synthesize underrepresented samples in training data for addressing the spurious correlation problem. To that end, we choose Stable Diffusion (SD) [9] for its established performance and accessibility. Our objective of addressing spurious correlation with SD necessitates a precise control over the (core/spurious) features of the generated content, which is not readily available in the original SD. Hence, we propose to first learn special token embeddings by tuning a text encoder to gain the ability of generating causal features for each class. Inspired by the textual inversion paradigm [6], we recreate samples of label $i$ using template sentences that incorporate a learnable token $[C_{i}]$ alongside the class name. For instance, in the Waterbirds dataset - a popular benchmark for the problem (see § 5.1), where the class name is "bird", our template sentence for the LandBird label would be:
@@
where $\tau_{\theta}(I)$ is the text encoder model, $I$ is the input prompt, $z_{t}$ is the latent noisy representation at time step $t$, $\epsilon$ is a noise sample, and $\epsilon_{\theta}$ is the denoising network. Note that, in the above, we also leverage the input image $x$ to ensure training data distribution aligned identification of the causal features. It is also noted that SD operates in the latent space of an auto-encoder. In Eq. (1), $\mathcal{E(.)}$ denotes the encoder of that model, which remains frozen in our technique. The learned token for the causal feature eventually gets incorporated into the input template sentences of our generative process to serve as a conditioning signal for SD - see Fig. 2 (a), which we use to alter causal parts of images.
@@ 5 Experiments
We evaluate our approach, on the standard vision benchmarks designed to assess spurious correlations and distribution shifts, comparing its performance against relevant baselines. Additionally, we conduct an ablation study to examine the impact of the pruning process and hyper-parameter selection.
@@ 5.1 Evaluation Settings
Compared Methods. We compared our method with the existing state-of-the-art methods that address spurious correlations, both with and without data augmentation. Specifically, we perform comparisons with DFR [12], JTT [15], MaskTune [25] and AFR [20] for methods without data augmentation. For methods incorporating data augmentation, we compared with DISC [28], LISA [30], and DAC [19]. For more details about these methods, see § 2.
Models. For a fair comparison, similar to other compared methods, we use a ResNet-50 model pre-trained on ImageNet as the classifier. In our approach, for textual inversion and sample generation, we leveraged , the Hugging Face implementation of Stable Diffusion [22] [^2] https://huggingface.co/stabilityai/stable-diffusion-2-inpainting. Additionally, for the pruning phase, we employed the PyTorch implementation of Integrated Gradients [24] [^3] https://pytorch.org/tutorials/beginner/introyt/captumyt.html.
Hyperparameters. For all datasets, hyperparameters were tuned to maximize worst-group accuracy on the validation set. We selected the lowest-loss $K$ portion of samples per class and the upweighting factor $\gamma$ from $\{0,0.1,0.2,\dots,1\}$ and $\{1,2,3,\dots,10\}$, respectively. The threshold $\mathrm{P}$ was determined using Integrated Gradients (IG) scores from a subset of generated samples.
@@ 5.2 Results
Among the methods that do not rely on data augmentation, our approach achieves superior worst-group accuracy—a key metric for group robustness—on Waterbirds and MetaShift. The MetaShift dataset, which involves domain shift, further highlights our method’s effectiveness in addressing such shifts by redirecting the ERM model’s attention to causal features. Unlike existing methods, our approach generates rare samples to enhance robustness, a capability absent in this category. On CelebA, our results remain competitive with state-of-the-art methods.
For methods leveraging data augmentation, our method also achieves superior worst-group accuracy on Waterbirds and MetaShift. LISA requires group labels for training, whereas our method operates without them while achieving comparable performance on CelebA. Similar to our approach, DaC addresses spurious correlations compositionally by combining causal and non-causal image components. However, DaC primarily mitigates spurious objects but remains vulnerable to spurious features. CelebA, which involves spurious feature correlations (e.g., hair length and label), presents a challenge for DaC, whereas the proposed method remains robust to both spurious objects and features, delivering competitive performance with state-of-the-art methods. For further details on different types of spurious correlations, we refer to the supplementary material.
@@
Dataset Number of samples Pruned NoPruning
@@ 5.3 Ablation Study
Pruning. As shown in Table 2, pruning operation enhances the performance of our approach. Inspecting the results, in the MetaShift dataset, $75\%$ of newly generated samples for the Cat class (i.e., Dog samples altered to Cat) are pruned. This is due to many Dog class images containing dogs that are either too small or entirely absent. Consequently, Stable Diffusion struggles to modify these regions effectively, leading to corrupted samples that require pruning. We discuss this more in the supplementary material.
@@
Similarly, a considerable number of the modified NotBlond images get pruned in Table 2. On inspection, we discover that many samples in this class feature bald individuals or those wearing hats, making it challenging for LangSAM to predict masks accurately. We also observed that LangSAM sometimes struggled with reliable hair region identification, leading to corrupted generated images. These results highlight the need of pruning in an automated pipeline of data augmentation to address the spurious correlation problem, and substantiate the efficacy of our method.
@@ 6 Conclusion
We introduced a novel approach for generating underrepresented samples in training data using text-to-image generative models. We leveraged diffusion models to selectively modify specific regions of images in a controlled manner to generate scarce samples, and mitigate spurious correlations. We employed textual inversion for this purpose. We also introduced a pruning mechanism that leverages attribution scores and ERM predictions to filter out undesired generative samples. Our results demonstrate that the proposed method effectively mitigates spurious correlations and enhances classification performance under distribution shifts. Although our primary exploration was about mitigating spurious correlations, which is shown by strong performance across a variety of datasets, our method has the potential to address general data scarcity problem.
Limitations and Ethical Considerations. Our method leverages existing diffusion and language-segmentation models. This potentially limits the applicability of our method as per the abilities of these models. However, with the rapid advances in generative and language models, this limitation is of minor nature. Continuously growing abilities of these models will proportionally strengthen our underlying framework. We also note that since our method provides precise automated control for image modification, it poses a potential risk of misuse in Deepfake generation. Hence, we advocate only a responsible use of our method for its intended purpose.
@@ 22             $f\leftarrow UpdateWeights(L_{total})$
Algorithm 2 T2I based balancing
@@ 8.2 Textual inversion
We selected minority group samples for the textual inversion dataset to mitigate the misperformance of the masking model. If the mask captures the spurious parts of the image, the Stable Diffusion model will also recreate these undesired features. For instance, in the CelebA dataset, where the mask is intended to isolate the hair, if the mask encompasses both the face and the hair, Stable Diffusion will regenerate the face in accordance with the learned representation of the minority group samples, provided the token has been trained to capture such features. As shown in Figure 5, the mask includes the face region as well. However, the learned token generates a non-blond woman, as the text embedding model was trained on a dataset of non-blond women.

### v2

# DDB: Diffusion Driven Balancing to Address Spurious Correlations

@@ Abstract
Deep neural networks trained with Empirical Risk Minimization (ERM) perform well when both training and test data come from the same domain, but they often fail to generalize to out-of-distribution samples. In image classification, these models may rely on spurious correlations that often exist between labels and irrelevant features of images, making predictions unreliable when those features do not exist. We propose a Diffusion Driven Balancing (DDB) technique to generate training samples with text-to-image diffusion models for addressing the spurious correlation problem. First, we compute the best describing token for the visual features pertaining to the causal components of samples by a textual inversion mechanism. Then, leveraging a language segmentation method and a diffusion model, we generate new samples by combining the causal component with the elements from other classes. We also meticulously prune the generated samples based on the prediction probabilities and attribution scores of the ERM model to ensure their correct composition for our objective. Finally, we retrain the ERM model on our augmented dataset. This process reduces the model’s reliance on spurious correlations by learning from carefully crafted samples in which this correlation does not exist. Our experiments show that across different benchmarks, our technique achieves better worst-group accuracy than the existing state-of-the-art methods. Our code is available at https://github.com/ArianYp/DDB.
@@ 1 Introduction
We refer to our method as Diffusion Driven Balancing (DDB). Inline with the compositional view of images [21], it leverages textual inversion [7] and identifies the optimal tokens for the causal parts of each class. It then employs language-based segmentation [15] to identify the visual causal counterparts in the images. These parts are later employed by our technique for diffusion-based image generation that can balance the data. In Fig. 1, we show representative examples of the generated minority group samples computed by our technique by altering the causal parts of the majority class images from the training data. A key challenge in generating such samples is ensuring that they truly capture the desired properties for addressing the spurious correlation problem. Otherwise, the impact of balancing can be detrimental. Our processing pipeline leverages multiple components, and under-performance by any of those can potentially lead to a low quality image. To handle that, we introduce an effective pruning strategy based on the attribution score and the prediction probability distribution of the ERM model. Our approach evaluates how much a generated sample influences the model’s predictions, ensuring that only samples with a strong causal impact are retained. Moreover, by further leveraging attribution scores, our approach ensures higher-quality relevant images, leading to improved group balancing and better generalization.
@@ 3 Problem setup
Given a dataset $D_{tr}=\{(x_{i},y_{i})\}_{i=0}^{N}$, a typical classifier $f$ is aimed at estimating $P(y|x)$. We model the problem with Structrul Causal We consider the compositional nature [33] of the samples $\{x_{i}\}_{i=0}^{N}$, where any sample consists of two types of features: core features ($c$) and spurious features ($s$). Core features (c) causally determine the correct label of the sample, while the spurious features (s) correlate with the label mainly due to irrelevant peculiarities of the dataset, e.g., class imbalances. These features are not inherently predictive. In other words, $P(y|s)\neq P(y)$.
@@
A dataset can be partitioned into majority and minority groups based on $(c,s)$ combinations, where majority samples contain a spurious feature ($s_{\text{maj}}$). In the Structural Causal Model (SCM) [22] view, this creates a backdoor path $s_{\text{maj}}\rightarrow Y$, leading the classifier to rely on those features. Consequently, it fails to generalize when the correlation breaks [32, 8]. Following common convention, we refer to the group on which $f$’s performance is worst as the worst group.
@@ 4 Method
We introduce DDB (Diffusion-Driven-Balancing), a three-stage approach designed to balance dataset groups effectively - see Fig. 2. Our first stage focuses on generating new data for minority groups using a diffusion-based generative model [25]. In the second stage, the generated data is pruned based on empirical risk minimization (ERM) loss and attribution scores, ensuring that only high-quality and relevant samples are retained. Finally, in the third stage, the ERM model is retrained on the newly balanced dataset, improving its robustness and mitigating spurious correlations. Below, we discuss these stages in detail.
Figure 2: An overview of DDB - Dog/Cat classes used for illustration. (a) First, the token $C_{dog}$ is optimized by reconstructing samples from the Dog class using prompts that include the trainable $C_{dog}$. Then, for each batch in $\mathcal{D}_{cat}$, masks $M$ are generated using the LangSAM [15] with the prompt Class Name = ‘animal’. These masks, along with the input images and a textual prompt, are fed into the inpainting model to generate new images. The token $C_{dog}$ is incorporated to generate the causal features of the Dog class. (b) The generated samples are then passed through the ERM model and the integrated gradients (IG) module to compute relevant scores for filtering undesired generated images. Algorithm for this stage is detailed in Alg. 1. (c) The final set of generated samples is used to retrain the ERM model.
@@ 4.1 Generating data
[…]
@@
where $\tau_{\theta}(I)$ is the text encoder model, $I$ is the input prompt, $z_{t}$ is the latent noisy representation at time step $t$, $\epsilon$ is a noise sample, $\epsilon_{\theta}$ is the denoising network, and input image $x$ from class $i$. It is also noted that SD operates in the latent space of an auto-encoder. In Eq. (1), $\mathcal{E(.)}$ denotes the encoder of that model, which remains frozen in our technique. The learned token for the causal feature eventually gets incorporated into the input template sentences of our generative process to serve as a conditioning signal for SD - see Fig. 2 (a), which we use to alter causal parts of images.
@@ 5 Experiments
We evaluate our approach, Diffusion-Driven Balancing (DDB), on the standard vision benchmarks designed to assess spurious correlations and distribution shifts, comparing its performance against relevant baselines. Additionally, we conduct an ablation study to examine the impact of the pruning process and hyper-parameter selection.
@@ 5.1 Evaluation Settings
Compared Methods. We compared DDB with the existing state-of-the-art methods that address spurious correlations, both with and without data augmentation. Specifically, we perform comparisons with DFR [14], JTT [17], MaskTune [28] and AFR [23] for methods without data augmentation. For methods incorporating data augmentation, we compared with DISC [31], LISA [33], and DAC [21]. For more details about these methods, see § 2.
Models. For a fair comparison, similar to other compared methods, we use a ResNet-50 model pre-trained on ImageNet as the classifier. In our approach, for textual inversion and sample generation, we leveraged the Hugging Face implementation of Stable Diffusion [25] [^2] https://huggingface.co/stabilityai/stable-diffusion-2-inpainting. Additionally, for the pruning phase, we employed the PyTorch implementation of Integrated Gradients [27] [^3] https://pytorch.org/tutorials/beginner/introyt/captumyt.html.
[…]
@@ 5.2 Results
Among the methods that do not rely on data augmentation, our approach achieves superior worst-group accuracy—a key metric for group robustness—on Waterbirds and MetaShift. The MetaShift dataset, which involves domain shift, further highlights DDB’s effectiveness in addressing such shifts by redirecting the ERM model’s attention to causal features. Unlike existing methods, our approach generates rare samples to enhance robustness, a capability absent in this category. On CelebA, our results remain competitive with state-of-the-art methods.
For methods leveraging data augmentation, DDB also achieves superior worst-group accuracy on Waterbirds and MetaShift. LISA requires group labels for training, whereas our method operates without them while achieving comparable performance on CelebA. Similar to our approach, DaC addresses spurious correlations compositionally by combining causal and non-causal image components. However, DaC primarily mitigates spurious objects but remains vulnerable to spurious features. CelebA, which involves spurious feature correlations (e.g., hair length and label), presents a challenge for DaC, whereas DDB remains robust to both spurious objects and features, delivering competitive performance with state-of-the-art methods. For further details on different types of spurious correlations, we refer to the supplementary material.
@@
Dataset Number of samples Pruned DDB - Pruning
@@ 5.3 Ablation Study
[…]
@@
Similarly, a considerable number of the modified NotBlond images get pruned in Table 2. On inspection, we discover that many samples in this class feature bald individuals or those wearing hats, making it challenging for LangSAM to predict masks accurately. We also observed that LangSAM sometimes struggled with reliable hair region identification, leading to corrupted generated images. These results highlight the need of pruning to substantiate the efficacy of our method. We provide more details in the supplementary material for failure types.
@@ 6 Conclusion
We introduced Diffusion Driven Balancing (DDB), a novel approach for generating underrepresented samples in training data using text-to-image generative models. We leveraged diffusion models to selectively modify specific regions of images in a controlled manner to generate scarce samples, and mitigate spurious correlations. We employed textual inversion for this purpose. We also introduced a pruning mechanism that leverages attribution scores and ERM predictions to filter out undesired generative samples. Our results demonstrate that the proposed method effectively mitigates spurious correlations and enhances classification performance under distribution shifts. Although our primary exploration was about mitigating spurious correlations, which is shown by strong performance across a variety of datasets, our method has the potential to address general data scarcity problem.
Limitations and Ethical Considerations. Our method leverages existing diffusion and language-segmentation models. This potentially limits the applicability of our method as per the abilities of these models. However, with the rapid advances in generative and language models, this limitation is of minor nature. Continuously growing abilities of these models will proportionally strengthen our underlying framework. We also note that since our method provides precise automated control for image modification, it poses a potential risk of misuse in Deepfake generation. Hence, we advocate only a responsible use of our method.
7 Acknowledgment
Naveed Akhtar is a recipient of the Australian Research Council Discovery Early Career Researcher Award (project # DE230101058) funded by the Australian Government. This work is also partially supported by Google Research Scholar Program Award.
@@ 22     $f\leftarrow UpdateWeights(L_{total})$
Algorithm 2 Diffusion-Driven-Balancing (DDB)
@@ 9.2 Textual inversion
We selected minority group samples for the textual inversion dataset to ensure that the learnable token does not reconstruct the spurious feature $s_{\text{maj}}$, thereby preventing error propagation when other components fail. For instance, in the CelebA dataset, where the mask is intended to isolate the hair, if the mask encompasses both the face and the hair, Stable Diffusion will regenerate the face in accordance with the learned representation of the minority group samples, provided the token has been trained to capture such features. As shown in Figure 5, the mask includes the face region as well. However, the learned token generates a non-blond woman, as the text embedding model was trained on a dataset of non-blond women.
@@ 9.3 Pruning
9.4 Imagenet-A
[…]
Figure 9: Examples of generated samples along with the corresponding masks and initial images. The generated images belong to the cockroach class, with initial images taken from the bee class, and vice versa.
9.5 Qualitative images
In this section, we provide additional qualitative examples across datasets. Figures 12 and 13 present generated images from the Waterbirds dataset, Figures 10 and 11 show results from the CelebA dataset, and Figures 14 and 15 show results from the Metashift dataset. Each figure showcases successful generations of minority group samples across different classes.
Figure 10: Examples of generated female samples along with the corresponding masks and initial images. The initial images are belong to male class.
Figure 11: Examples of generated male samples along with the corresponding masks and initial images. The initial images are belong to female class.
Figure 12: Examples of generated landbird samples along with the corresponding masks and initial images. The initial images are belong to waterbird class.
Figure 13: Examples of generated waterbird samples along with the corresponding masks and initial images. The initial images are belong to landbird class.
Figure 14: Examples of generated dog samples along with the corresponding masks and initial images. The initial images are belong to cat class.
Figure 15: Examples of generated cat samples along with the corresponding masks and initial images. The initial images are belong to dog class.